Outdoor Outfitters has created a flexible budget for the 70,000-unit and the 80,000-unit levels of activity shown as follows. Complete Outdoor Outfitters’s flexible budget at the 90,000-unit level of activity. Assume that the cost of goods sold and variable operating expenses vary directly with sales and that income taxes remain at 30 percent of operating income.

70,000 Units 80,000 Units 90,000 Units
Sales $1,400,000 $1,600,000
Cost of goods sold 840,000 960,000
Gross profit on sales $560,000 $640,000 $0
Operating expenses ($90,000 fixed) 370,000 410,000
Operating income $190,000 $230,000 $0
Income taxes (30% of operating income) 57,000 69,000
Net income $133,000 $161,000 $0
Answer #1
For For
70000 units Per unit 80000 units Per unit
Sales Per unit 1400000/70000 20 1600000/80000 20
Cost per unit 840000/70000 12 960000/80000 12

Calculation of operating expenses which is fixed

For 70000 units

90000 is fixed

so from total operating expenses we will deduct fixed cost and then divide it by no of units

370000 - 90000 = 280000

For Remaining 280000 / 70000 = 4 per unit

For 80000 units

90000 is fixed

so from total operating expenses we will deduct fixed cost and then divide it by no of units

410000 - 90000 = 320000

For Remaining 320000 / 80000 = 4 per unit

Flexible budget for 90000 units

90000 units
A Sales 90000 x 20 1800000
B Cost of goods sold 90000x12 1080000
C Gross profit on sales A-B 720000
D operating expenses 90000 + (4 X 90000) 450000
E operating income C-D 270000
F Income Taxes E X 30% 81000
G Net Income E-F 189000

Sales will be calculated by multiplying 90000 units x sale per unit (calculated from 70000 and 80000 units)

and Cost of goods sold will be calculated by multiplying 90000 units x cost per unit (calculated from 70000 and 80000 units)

Operating Expenses = 90000 will be fixed

& 90000 x 4 = 360000

Total 450000

Gross profit = sales -Cost of goods sold

operating income = Gross profit - Operating expenses

Income Taxes = operating income x 30%

Net Income = operating income -Income Taxes